OpenStudy (anonymous):

Solve the equation. W/9 = 9.....do u divide this?

OpenStudy (anonymous):

jjbaez go and sleep

OpenStudy (anonymous):

i need actually lol

OpenStudy (anonymous):

If W/9=something then W=9 times something.

OpenStudy (anonymous):

B / 2 - 15 = 16 i dont get how to do equations like this

OpenStudy (anonymous):

Multiply every term by 2 so there are no fractions and add 30 to both sides so you get B=something

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

see u need to keep the variables on one side and the values on the other side like in this case u will need to keep B on one side and values on the other side like this one B/2 = 16 + 15

OpenStudy (anonymous):

B/2 = 31

OpenStudy (anonymous):

therefore cross multiply it !

OpenStudy (anonymous):

ok i c